Chapter 03, Problem 017 rhree vectors a, b, and T, each have a magnitude of 45.0 m and lie in an xy plane. Their directions relative to the positive direction of the x axis are 25.0 (-180° 180°) and (c) the magnitude and (d) the angle of a b + ·in the range of (-180e, i 80e)? what are (e) the magnitude and (f) the angle (in the enaciean T 192 , and 313,respectively. What are (a) the magnitude and (b) the angle of the vector (relative to the tx direction in the range of range of (-1809, 180°)) of a fourth vector d such that a-+d) o?
Answer #1

a)

overrightarrow{a}=45Cos25widehat{i}+45Sin25widehat{j}

b45Cos192i +45Sin 192j

overrightarrow{c}=45Cos313widehat{i}+45Sin313widehat{j}

ที่ + + -(27.457)it (-23.249)

Magnitude

ス+ b·ス| = V/27.4572 + (-23.249)-35.98m

b)

-23.249 27.457 -40.26° θ tan-1

c)

スーb +ス= 1 15.49i + (-4.537)j

Magnitude

left | overrightarrow{a}-overrightarrow{b}+overrightarrow{c} ight |=sqrt{115.49^{2}+(-4.537)^{2}}=115.58 m

d)

direction

/一4.537 115.49 θ = tan-1 =-2.25°

e)

(overrightarrow{a}+overrightarrow{b})-left ( overrightarrow{c}+overrightarrow{d} ight )=0

6

d(-33.923)i 42.573j m

Magnitude

(-33.923)(42.573)2 54.435m

f)

heta =tan^{-1}left ( rac{42.573}{-33.923} ight )=180-51.45=128.55^{o}